OpenStudy (anonymous):

Express answer in exact form. Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.)

OpenStudy (anonymous):

For an inscribed square in a circle, the diagonal of the square is equal to the diameter of the square. So the diagonal of the square= sqrt(2)*6 Diameter=sqrt(2)*6 Radius=sqrt(2)*3 Now the square divides the circle into 4 sectors of equal area. So just subtract the area of the square from that of the circle. This gives pi*(sqrt(2)*3)^2-6*6 =18pi-36 Now this is the area of all four segments,therefore the area of one segment is (18pi-36)/4 =4.5pi-9